OpenStudy (anonymous):

Find the surface area of a sphere with a diameter of 14 cm. Leave your answer in terms of pi @Directrix

Directrix (directrix):

A = 4 * π * r² A = 4 * π * 7 ²

Directrix (directrix):

The diameter is twice the radius so you have to take half the diameter of 14 to get the radius.

OpenStudy (anonymous):

i got 615.75 ..did i do it wrong

Directrix (directrix):

The instructions say to give the answer in terms of pi. That means do not put 3.14 for pi.

Directrix (directrix):

I got 196π

OpenStudy (anonymous):

sorry i did that

OpenStudy (anonymous):

thats it

Directrix (directrix):

next question